18. 2h^2-h-3=0
2(h-1.5)(h+1)
Tried to work it out don't know if this is the right answer I do not get how to doe these. Can you please check thank you very much I don't know what I would do with out such good help like you guy and gals. Thanks again

Answer by stanbon(75887)

2h^2-h-3=0
2h^2-3h+2h-3=0
h(2h-3)+(2h-3)=0
(2h-3)(h+1)=0
h=3/2 or h=-1
Cheers,
Stan H.